# CLIMB-3D: Continual Learning for Imbalanced 3D Instance Segmentation

### v1

@@ 4.1 Problem Formulation
The objective of 3D point cloud instance segmentation is to accurately identify and segment individual instances of objects within a given point cloud. Mathematically, the training dataset is represented as $\mathcal{D}=(\mathcal{P},\mathcal{Y})=\{(p_{i},y_{i})\}_{i=1}^{N}$, where $N$ is the total number of samples. Each sample consists of a colored point cloud $p_{i}\in\mathbb{R}^{M\times 6}$ of size $M$, where the point coordinates and color values are represented as $\{x,y,z,r,g,b\}$. The corresponding annotations are denoted as $y_{i}=\{(m_{i,j},c_{i,j})\}_{j=1}^{J}$, where $m_{i,j}$ represents the instance mask for the $j$-th instance, and $c_{i,j}\in\mathcal{C}=\{1,\dots,C\}$ denotes the semantic label of the object category to which the instance belongs for the $i$-th point cloud. Here, $J$ represents the total number of instances in the $i$-th point cloud, and $C$ indicates the number of distinct object categories. During the learning process, the model $\Phi$ will process this dataset and output predictions $\hat{y}_{i,j}=(\hat{m}_{i,j},\hat{c}_{i,j})$, where $\hat{m}_{i,j}$ represents the predicted instance mask and $\hat{c}_{i,j}$ denotes the predicted semantic label for the $j$-th instance in the $i$-th point cloud.
To adapt the dataset to an incremental learning setting, we partition the object categories $\mathcal{C}$ into $T$ subsets, denoted as $\mathcal{C}=\mathcal{C}^{1}\cup\cdots\cup\mathcal{C}^{T}$. Each phase $t\in\{1,\dots,T\}$ is associated with a specific subset $\mathcal{C}^{t}$, and its corresponding dataset $\mathcal{D}^{t}$ which only contains annotations for objects belonging to the corresponding subset. Formally, during the $t$-th phase of training, the dataset $\mathcal{D}^{t}=(\mathcal{P},\mathcal{Y}^{t})=\{(p_{i},y_{i}^{t})\}_{i=1}^{N}$ is defined, where $\mathcal{P}$ represents the point clouds shared across all phases, and $\mathcal{Y}^{t}$ contains annotations exclusively for objects belonging to the class subset $\mathcal{C}^{t}$. It is important to note that the 3D scenes within each phase can contain objects of any type from the entire object category set $\mathcal{C}$, but only the object belonging to $\mathcal{C}^{t}$ are annotated during that specific phase. After training for phase $t$ completes, the model is evaluated on a validation set containing the union of classes up to task $t$ (i.e., $\mathcal{C}^{1}\cup\cdots\cup\mathcal{C}^{t}$). Training progresses to the next phase, $t+1$, where the model $\Phi$ observes the same set of 3D scenes $\mathcal{P}$ but with annotations for different object types belonging to the subset $\mathcal{C}^{t+1}$. This incremental training approach allows the model to gradually learn and adapt to new object categories over multiple phases.
4.2 CLIMB-3D
In our proposed framework (Figure 2), the incremental instance segmentation model undergoes phased training, as described in Sec. 4.1, where carefully designed subsets of the dataset are introduced to handle various real-world scenarios discussed in Sec. 4.3. Formally, at phase $t$, when the model $\Phi^{t}$ is introduced with input data $\mathcal{D}^{t}=\{(p_{i},y_{i}^{t})\}_{i=1}^{N}$ and trained using Eq. 1, a common issue arises where it tends to forget the knowledge acquired in the previous phase, leading to catastrophic forgetting [mccloskey1989catastrophic]. To address this, we first draw inspiration from techniques developed in the 2D domain [rebuffi2017icarl, li2017learning] and recent 3D semantic segmentation [su2024balanced, Yang_2023_CVPR], and adapt them for our setting. However, we observe that these adaptations alone fall short of achieving the desired performance levels; therefore, we propose a teacher-student knowledge distillation framework to effectively retain previously learned knowledge. Additionally, we incorporate an imbalance correction module to handle the challenge of less frequent classes from earlier tasks.
Exemplar Replay (ER). Inspired by the approach proposed by Buzzega et al. [buzzega2020dark], ER methods alleviate the issue of limited exposure to previous task data during training. By selectively storing a small subset of exemplars $\mathcal{E}_{t}$ from previous phases, the model can learn from both the current task data $\mathcal{D}_{t}$ and the replayed exemplars $\mathcal{E}_{1:t-1}$. This results in a combined dataset $\mathcal{D}_{t}\cup\mathcal{E}_{1:t-1}$, where $\mathcal{E}_{1:t-1}$ represents the exemplar memory formed by the union of all previous exemplar sets $\mathcal{E}_{1:t-1}=\mathcal{E}_{1}\cup~{}\dots~{}\cup\mathcal{E}_{t-1}$. The model undergoes a full iteration on $\mathcal{D}_{t}$ before replaying the exemplars.
Some previous 3D approaches adopt this strategy to retain knowledge but rely on a large exemplar set [boudjoghra20243d], which is often impractical in real-world scenarios. To address this, we choose a smaller exemplar set, creating a more challenging setup that requires the model to effectively manage and retain knowledge with limited resources, thereby testing its robustness and adaptability in practical applications.
Knowledge Distillation (KD) Module. In our incremental learning approach, we utilize a Knowledge Distillation (KD) module that incorporates a teacher-student framework, maintaining a copy of the previously trained model while learning the current task. For $t>1$, at the beginning of each training stage, the current model $\Phi^{t}$ is initialized as $\Phi^{t}\leftarrow\Phi^{t-1}$, where $\Phi^{t-1}$ represents the model trained in the previous phase. As the $\Phi^{t}$ is trained on the previous stage dataset, it holds information about the previous set of classes. Hence, we make use of this model to help retain previous knowledge while learning the current task.
When presented with a new training point-cloud and label pair $(p,y^{t})$, the output of the previous model is calculated as $\hat{y}_{\mathrm{pseudo}}=\Phi^{\mathrm{t-1}}(p)$, and a combined loss function is minimized. This combined loss comprises the $\mathcal{L}_{\mathrm{Seg}}(\hat{y}^{t},y^{t})$ loss, which measures the discrepancy between the predicted and ground truth labels, and the knowledge distillation loss $\mathcal{L}_{\mathrm{KD}}$, which encourages the similarity between the predictions of the current model and the previous model.
$\mathcal{L}_{\mathrm{KD}}(\hat{y}^{t},\hat{y}_{\mathrm{pseudo}})$ $=\mathcal{L}_{\mathrm{mask}}(\hat{m}^{t}_{j},\hat{m}_{j,\mathrm{pseudo}})$
$\quad\quad\quad\quad+\lambda_{\mathrm{cls}}\mathcal{L}_{\mathrm{cls}}(\hat{c}^{t}_{j},\hat{c}_{j,\mathrm{pseudo}})$ (2)
However, as pointed out in previous works on object detection in 2D [liu2023continual], Section 4.2 is often biased towards the background classes, as the model tends to predict the background for most instances. Similarly, we propose selecting the top $K$ most confident predictions from the previous teacher model $\Phi^{t-1}$ and combining them with the ground truth labels, which then serve as pseudo-labels. By extracting the top $K$ confident samples from the output of $\Phi^{t-1}$ and combining them with the ground truth labels, the augmented label set becomes $y^{\prime}=[y^{t},\hat{y}_{\mathrm{pseudo}}^{K}]$. This augmented label set is then used to optimize the current model $\Phi$ with the $\mathcal{L}_{\mathrm{Seg}}$ loss function from Equation 1.
Imbalance Correction (IC) Module. Although retaining a few samples from previous tasks and selecting the most confident predictions from the previous model, $\Phi^{t-1}$, helps preserve information from prior tasks while learning new ones, we observed that this approach does not adequately address the class imbalance. Our analysis reveals that the most confident predictions from $\Phi^{t-1}$ are largely associated with the most frequent object categories, causing the model to forget less common classes. This issue can be mitigated by re-weighting the predictions based on the frequency of observed categories [islam2021class].
At task $t$, we only have access to the data and statistics of the current task; the previous task’s dataset and statistics are unavailable. To incorporate balancing elements despite the absence of previous stage statistics, we propose leveraging pseudo prediction statistics from $\Phi^{t-1}$. During each iteration, we use $\Phi^{t-1}$ to generate pseudo labels and accumulate class frequency statistics for prior tasks throughout the current epoch training. At the end of each epoch, we combine the statistics of observed scene classes and predicted pseudo-classes, calculating the frequency $\mathbf{f}$ of all classes seen so far.
Formally, for each category $c$, we assign a weight $\mathbf{w}_{c}$ inversely proportional to its observed frequency in the predictions of $\Phi^{t-1}$ and the current dataset. The weight $\mathbf{w}_{c}$ is defined as: $\mathbf{w}_{c}=\frac{1}{\mathbf{f}(c)+\epsilon}$, where $\epsilon$ is a small constant to avoid division by zero. In the next epoch, predictions from $\Phi^{t-1}$ are re-weighted using $\mathbf{w}_{c}$, creating an adjusted high-confidence pseudo label set for less frequent categories: $\hat{y}_{\mathrm{pseudo}}^{(w)}=\mathbf{w}_{c}\cdot\hat{y}_{\mathrm{pseudo}}$. This re-weighting occurs at each epoch, allowing $\Phi^{t-1}$ to yield a broader set of less frequent classes. To ensure the model encounters both high-confidence and less frequent classes, we select the top $K$ high-confidence predictions both before and after re-weighting. The resulting augmented label space, which combines ground truth labels, original pseudo labels, and re-weighted pseudo labels, is given by: $y^{\prime}=[y,\hat{y}_{\mathrm{pseudo}}^{K},y_{\mathrm{pseudo}}^{(w),K}]$.
As this weighting scheme is applied only to previous model predictions, we further tune the current model to favor less frequent classes by incorporating the same weights $\mathbf{w}_{c}$ into the classification loss of Equation 1. The adjusted segmentation loss becomes:
$\mathcal{L}_{\mathrm{3DIS}}(y^{\prime}_{j},\hat{y}_{j})=\mathcal{L}_{\mathrm{mask}}(m^{\prime}_{j},\hat{m}_{j})+\mathbf{w}^{\prime}_{c}\mathcal{L}_{\mathrm{cls}}(c^{\prime}_{j},\hat{c}_{j}),$ (3)
where $\mathbf{w}^{\prime}_{c}=\mathbf{w}_{c}\cdot\lambda_{\mathrm{cls}}$ represents the adjusted category weights. By re-weighting and augmenting both the label space and the loss function, our IC module ensures that both the current and previous models encounter pseudo labels spanning the long-tail distribution, addressing class imbalance and enabling more balanced learning.
4.3 Designing Incremental Scenarios
Figure 3: Incremental scenarios are grouped based on frequency of occurrence, semantic similarity, and random clustering. Different color clouds ( ,  ,  ) represent tasks in each scenario, while various shapes represent object categories, and   denotes the background. Left: Tasks are organized based on the frequency of object categories. Middle: Tasks are grouped by semantic similarity, where objects with similar shapes (e.g., circles, plus signs, and triangles) denote semantically similar classes. Right: In this fully random scenario, tasks may contain a mix of semantically similar, more frequent, or less frequent classes.
While conventional incremental learning methods have numerous practical applications, they often assume an equal distribution of samples, which does not reflect real-world conditions. In practice, the number of object categories, $\mathcal{C}$, is typically large, with significant variability in category occurrence, shape, structure, and size. With these attributes in mind, we design three incremental learning scenarios, each addressing distinct aspects of real-world conditions, the design is highlighted in Figure 3.
1

### v2

@@ 3.1 Problem Formulation
3DIS. The objective of this task is to detect and segment each object instance within a given point cloud. Formally, the training dataset is denoted by $\mathcal{D}=\{(\mathbf{P}_{i},\mathbf{Y}_{i})\}_{i=1}^{N}$, where $N$ is the number of training samples. Each input $\mathbf{P}_{i}\in\mathbb{R}^{M\times 6}$ is a colored point cloud consisting of $M$ points, where each point is represented by its 3D coordinates and RGB values. The corresponding annotation $\mathbf{Y}_{i}=\{(m_{i,j},c_{i,j})\}_{j=1}^{J_{i}}$ contains $J_{i}$ object instances, where $m_{i,j}\in\{0,1\}^{M}$ is a binary mask indicating the points belonging to the $j$-th instance, and $c_{i,j}\in\mathcal{C}=\{1,\dots,C\}$ is the semantic class label of that instance. Given $\mathbf{P}_{i}$, the model $\Phi$ predicts instance-level outputs $\hat{\mathbf{Y}}_{i}=\{(\hat{m}_{i,j},\hat{c}_{i,j})\}_{j=1}^{\hat{J}_{i}}$, where $\hat{m}_{i,j}$ and $\hat{c}_{i,j}$ denote the predicted mask and category label for the $j$-th instance, respectively. The number of predicted instances $\hat{J}_{i}$ varies depending on the model’s inference. The model is optimized using below objective:
$\mathcal{L}_{\text{3DIS}}(\mathcal{D};\Phi)$ $=\frac{1}{|\mathcal{D}|}\sum_{(\mathbf{P},\mathbf{Y})\in\mathcal{D}}\frac{1}{|\mathbf{Y}|}\sum_{(m_{j},c_{j})\in\mathbf{Y}}$
$\quad\left(\mathcal{L}_{\text{mask}}(m_{j},\hat{m}_{j})+\lambda\cdot\mathcal{L}_{\text{cls}}(c_{j},\hat{c}_{j})\right)$ (1)
where $\mathcal{L}_{\text{cls}}$ and $\mathcal{L}_{\text{mask}}$ are the average classification and mask losses over all instances, and $\lambda$ controls their trade-off.
CI-3DIS. Unlike conventional 3DIS, CI-3DIS involves sequential learning, where object categories are introduced incrementally over $T$ training tasks. Each task $t$ introduces a disjoint set of classes $\mathcal{C}^{t}$, with $\mathcal{C}=\bigcup_{t=1}^{T}\mathcal{C}^{t}$ and $\mathcal{C}^{t}\cap\mathcal{C}^{t^{\prime}}=\emptyset$ for $t\neq t^{\prime}$. At each task $t$, the model receives a dataset $\mathcal{D}^{t}=\{(\mathbf{P}_{i},\mathbf{Y}_{i}^{t})\}_{i=1}^{N}$, where each coloured point cloud $\mathbf{P}_{i}\in\mathbb{R}^{M\times 6}$ contains $M$ points, and the corresponding annotation $\mathbf{Y}_{i}^{t}=\{(m_{i,j}^{t},c_{i,j}^{t})\}_{j=1}^{J_{i}^{t}}$ includes instance masks $m_{i,j}^{t}\in\{0,1\}^{M}$ and semantic labels $c_{i,j}^{t}\in\mathcal{C}^{t}$. The model $\Phi^{t}$, initialized from $\Phi^{t-1}$, is trained on $\mathcal{D}^{t}$ to predict instance-level outputs $\hat{\mathbf{Y}}_{i}^{t}=\{(\hat{m}_{i,j}^{t},\hat{c}_{i,j}^{t})\}_{j=1}^{\hat{J}_{i}^{t}}$, where each predicted class label $\hat{c}_{i,j}^{t}$ belongs to the cumulative set of all classes observed so far: $\bigcup_{k=1}^{t}\mathcal{C}^{k}$. The optimisation objective for the CI-3DIS setting is defined as:
$\mathcal{L}_{\text{CI-3DIS}}(\mathcal{D}^{t};\Phi^{t})$ $=\frac{1}{|\mathcal{D}^{t}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in\mathcal{D}^{t}}\frac{1}{|\mathbf{Y}|}\sum_{(m^{t}_{j},c^{t}_{j})\in\mathbf{Y}^{t}}$
$\quad\left(\mathcal{L}_{\text{mask}}(m^{t}_{j},\hat{m}^{t}_{j})+\lambda\cdot\mathcal{L}_{\text{cls}}(c^{t}_{j},\hat{c}^{t}_{j})\right).$ (2)
The key challenge in this setup is to learn new classes without forgetting those encountered in previous tasks, despite limited supervision and the absence of past labels.
3.2 Method: CLIMB-3D
Overview. As illustrated in Fig. 2 and introduced in Sec. 3.1, the proposed framework for CI-3DIS follows a phase-wise training strategy. In each phase, the model is exposed to a carefully selected subset of the dataset, designed to simulate real-world scenarios discussed in Sec. 3.3. Naïvely training on such phased data leads to catastrophic forgetting [mccloskey1989catastrophic], where the model forgets previously acquired knowledge when learning new tasks. To address this, we first incorporate Exemplar Replay (ER) [buzzega2020dark], inspired by 2D [rebuffi2017icarl, li2017learning, aljundi2018memory] and 3D [boudjoghra20243d] incremental setting, to mitigate forgetting by storing a small set of representative samples from earlier phases. However, ER alone is insufficient for achieving robust performance. We therefore introduce two additional components: a Pseudo-Label Generator (PLG), which leverages a frozen model to generate supervision signals for previously seen classes, and a Class-Balanced Re-weighting (CBR) module that compensates for class imbalance across phases. Each of these components is described in detail below  [^2] For clarity, modifications introduced by each component – ER, PLG, and CBR – are highlighted in green, blue, and red, respectively..
Exemplar Replay (ER). Inspired by Buzzega et al. [buzzega2020dark], ER addresses the issue of catastrophic forgetting by allowing the model to retain a small portion of data from earlier phases. At phase $t$, the model is trained not only on the current task data $\mathcal{D}^{t}$ but also on a set of exemplars $\mathcal{E}^{1:t-1}$ collected from previous phases. These exemplars are accumulated incrementally, where $\mathcal{E}^{1:t-1}=\mathcal{E}^{1}\cup\dots\cup\mathcal{E}^{t-1},$ and $\mathcal{E}^{t}$ denotes the exemplar subset stored after phase $t$, with $|\mathcal{E}^{t}|$ as the exemplar replay size. We define the combined replay dataset as: ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}=\mathcal{D}^{t}\cup\mathcal{E}^{1:t-1}.$ Epoch training proceeds in two stages: first on the current data $\mathcal{D}^{t}$, followed by replay using $\mathcal{E}^{1:t-1}$. The updated objective function for CI-3DIS, based on LABEL:eq:ci3dis-loss, becomes:
$\mathcal{L}_{\text{CI-3DIS}}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})$ $=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|\mathbf{Y}^{t}|}\sum_{(m_{j}^{t},c_{j}^{t})\in\mathbf{Y}^{t}}$
$\quad\left(\mathcal{L}_{\text{mask}}(m_{j}^{t},\hat{m}_{j}^{t})+\lambda\cdot\mathcal{L}_{\text{cls}}(c_{j}^{t},\hat{c}_{j}^{t})\right).$ (3)
While previous CI-3DIS approaches use large exemplar sets to mitigate forgetting [boudjoghra20243d], such strategies become impractical under memory constraints. Consequently, although exemplar replay (ER) aids knowledge retention, it alone proves insufficient for robust performance with a limited exemplar budget. We therefore propose PLG and CBM to better preserve past representations and enhance generalisation across phases.
Pseudo-Label Generator (PLG). In the CI-3DIS setting, at phase $t$ (for $t>1$), although we no longer have access to the ground-truth labels for previously seen classes, the model from the previous phase, $\Phi^{t-1}$, remains available and retains knowledge acquired from past tasks. We use this model to generate pseudo-labels for previous classes, thereby providing approximate past supervision during training. This allows the current model $\Phi^{t}$ to retain past knowledge and reduces forgetting, even without access to original annotations.
Given a point cloud and label pair $(\mathbf{P},\mathbf{Y}^{t})$ from $\mathcal{D}^{t}$, the previous model model $\Phi^{t-1}$ generates pseudo-labels for the previously learned classes as $\hat{\mathbf{Y}}^{1:t-1}=\Phi^{t-1}(\mathbf{P})$, which we denote as $\tilde{\mathbf{Y}}^{t}$ for brevity. However, we observe that using all predictions from $\Phi^{t-1}$ introduces noisy or incorrect labels, which degrade overall performance. To mitigate this, we select the top-$K$ most confident instance predictions, denoted as $\tilde{\mathbf{Y}}^{t}_{K}$, and concatenate them with the ground-truth labels of the current task to obtain the supervision set: ${\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{\mathbf{Y}}^{t}}=\mathbf{Y}^{t}\mathbin{\|}{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\tilde{\mathbf{Y}}^{t}_{K}}$. At the end, Using this supervision set, the model is optimized via the following updated objective:
$\mathcal{L}_{\text{CI-3DIS}}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})$ $=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{\mathbf{Y}}^{t}}|}\sum_{({\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{m}_{j}^{t}},{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{c}_{j}^{t}})\in{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{\mathbf{Y}}^{t}}}$
$\quad\left(\mathcal{L}_{\text{mask}}({\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{m}_{j}^{t}},\hat{m}_{j}^{t})+\lambda\cdot\mathcal{L}_{\text{cls}}({\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{c}_{j}^{t}},\hat{c}_{j}^{t})\right),$ (4)
where the loss is computed over both the current task’s ground-truth labels and the top-$K$ confident pseudo-labels from previous phases, enabling the model to retain prior knowledge while adapting to new classes.
Class-Balanced Re-weighting (CBR). While ER and PLG contribute to preserving prior knowledge, we observe that the model’s predictions are biased toward frequently occurring classes, leading to forgetting of rare categories. This issue is further reinforced by the top-$K$ pseudo-labels selected by PLG, which predominantly represent dominant classes in the data. To alleviate this issue, we introduce a Class-Balanced Re-weighting (CBR) scheme that adjusts for class imbalance based on frequency of object occurrence. However, at task $t$, we only have access to the dataset $\mathcal{D}^{t}$ and its corresponding label distribution $p^{t}(c)$ over classes $c\in\mathcal{C}^{t}$. The datasets and class distributions from previous tasks $\{\mathcal{D}^{i},p^{i}(c)\}_{i=1}^{t-1}$ are no longer accessible – making it challenging to directly account for class imbalance across all seen tasks. Therefore, we propose to leverage the pseudo-label predictions of the frozen model $\Phi^{t-1}$ as a proxy for the class distribution across previously learned categories.
During each training iteration, we apply the previous model $\Phi^{t-1}$ to the current input point cloud $\mathbf{P}$ to generate pseudo-labels $\tilde{\mathbf{Y}}^{t}$ , from which we accumulate class-wise frequency statistics corresponding to prior tasks. Let $\mathcal{C}^{1:t}=\bigcup_{i=1}^{t}\mathcal{C}^{i}$ represent the union of all classes seen up to task $t$. At the end of each epoch, we compute the overall class frequency $\mathbf{f}\in\mathbb{R}^{|\mathcal{C}^{1:t}|}$ by combining the pseudo-label distribution $\tilde{p}^{t}(c)$ derived from $\tilde{\mathbf{Y}}^{t}$ with the observed ground-truth label distribution $p^{t}(c)$ from the current task. The resulting frequency vector $\mathbf{f}=[f(c)]_{c\in\mathcal{C}^{1:t}}$ is then used to compute class-wise $\mathbf{w}=[w(c)]_{c\in\mathcal{C}^{1:t}}$ weights for balanced pseudo-label selection and to tune current model to favour rare classes.
During the next epoch, the model $\Phi^{t-1}$ produces soft pseudo-label predictions $\tilde{\mathbf{Y}}^{t}$ , where each component $\tilde{\mathbf{Y}}^{t}[c]$ denotes the confidence score for class $c\in\mathcal{C}^{1:t-1}$. To promote balanced pseudo-label selection, we apply class-wise re-weighting using $w(c)=\frac{1}{f(c)+\epsilon}$, where $\epsilon$ is a small constant for numerical stability. The adjusted predictions are computed by applying the class-wise weights as: $\tilde{\mathbf{Y}}^{t,w}[c]=w(c)\cdot\tilde{\mathbf{Y}}^{t}[c],\quad\forall c\in\mathcal{C}^{1:t-1}.$ We then perform top-$K$ selection over both the original and the re-weighted scores, yielding two pseudo-label sets: $\tilde{\mathbf{Y}}^{t}_{K}$ and $\tilde{\mathbf{Y}}^{t,w}_{K}$, respectively. The final supervision label set $\overline{\mathbf{Y}}^{t}$ is constructed by concatenating the ground-truth labels $\mathbf{Y}^{t}$ with both sets of pseudo-labels: ${\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}=\mathbf{Y}^{t}\mathbin{\|}{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\tilde{\mathbf{Y}}^{t}_{K}}\mathbin{\|}{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\tilde{\mathbf{Y}}^{t,w}_{K}}.$
Although the above re-weighting mitigates bias from prior tasks, it does not address class imbalance within the current task $t$. To resolve this, we extend the class-balanced weights $w_{c}$ to incorporate statistics from both previously seen and current classes, i.e., over $\mathcal{C}^{1:t}$ rather than only $\mathcal{C}^{1:t-1}$, and denote the updated weights as $w_{c}^{\prime}$, which are then integrated into the training objective. After incorporating all components, the modified objective in Eq. 1 becomes the final updated objective used in CLIMB-3D:
$\mathcal{L}_{\text{CI-3DIS}}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})$ $=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}|}\sum_{({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{m}_{j}^{t}},{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{c}_{j}^{t}})\in{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}}$
$\quad\left(\mathcal{L}_{\text{mask}}({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{m}_{j}^{t}},\hat{m}_{j}^{t})+{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime\prime}_{c}}\cdot\mathcal{L}_{\text{cls}}({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{c}_{j}^{t}},\hat{c}_{j}^{t})\right),$ (5)
where ${\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime\prime}_{c}}={\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime}_{c}}\cdot\lambda_{\mathrm{cls}}$ denotes the scaled weight vector, and $\lambda_{\mathrm{cls}}$ is a loss-balancing hyperparameter. To address the class-imbalance across tasks, our CBR module jointly re-weights pseudo-label selection and classification loss, ensuring balanced supervision across all seen categories, including rare ones. This unified strategy reduces bias and enhances performance, particularly for under-represented classes.
3.3 Benchmarking Incremental Scenarios
While CI-3DIS methods offer a wide range of practical applications, they frequently rely on the assumption of uniform sample distribution – an assumption that rarely holds in real-world settings. In practice, the number of object categories, denoted by $\mathcal{C}$, is often large and characterised by substantial variability in category frequency, shape, structure, and size. To address these challenges, we propose three incremental learning scenarios, each designed to capture a different facet of real-world complexity. The design of these scenarios is detailed below; for more information and illustration, please refer Appendix Appendix C of the supplementary material.